Financial statements for Remington Inc. follow.

Remington Inc.
Consolidated Statements of Income
(In thousands except per share amounts)
2019 2018 2017
Net sales $7,245,088 $6,944,296 $6,149,218
Cost of goods sold (5,286,253) (4,953,556) (4,355,675)
Gross margin $1,958,835 $1,990,740 $1,793,543
General and administrative expenses (1,259,896) (1,202,042) (1,080,843)
Special and nonrecurring items 2,617 - -
Operating income $701,556 $788,698 $712,700
Interest expense (63,685) (62,398) (63,927)
Other income 7,308 10,080 11,529
Gain on sale of investments - 9,117 -
Income before income taxes $645,179 $745,497 $660,302
Provision for income taxes (254,000) (290,000) (257,000)
Net income $391,179 $455,497 $403,302
Net income per share $1.08 $1.25 $1.11
Remington Inc.
Consolidated Balance Sheets
(In thousands)
ASSETS Dec. 31, 2019 Dec. 31, 2018
Current assets:
Cash and equivalents $320,558 $41,235
Accounts receivable 1,056,911 837,377
Inventories 733,700 803,707
Other 109,456 101,811
Total current assets $2,220,625 $1,784,130
Property and equipment, net 1,666,588 1,813,948
Other assets 233,792 248,372
Total assets $4,121,005 $3,846,450
LIABILITIES AND STOCKHOLDERS' EQUITY
Current liabilities:
Accounts payable $250,363 $309,092
Accrued expenses 347,892 274,220
Other current liabilities 15,700 -
Income taxes 93,489 137,466
Total current liabilities $707,444 $720,778
Long-term debt $650,000 $541,639
Deferred income taxes 275,101 274,844
Other long-term liabilities 61,267 41,572
Total liabilities $1,693,812 $1,578,833
Stockholders' equity:
Common and preferred stock $189,727 $189,727
Additional paid-in capital 114,806 127,776
Retained earnings 2,397,112 2,136,794
$2,701,645 $2,454,297
Less: Treasury stock, at cost (274,452) (186,680)
Total stockholders' equity $2,427,193 $2,267,617
Total liabilities and stockholders' equity $4,121,005 $3,846,450

Use the information above and below to respond to the following requirements.

Statement Item January 1, 2018 (In Millions)
Total assets $3,485,233
Total stockholders' equity   2,083,122
Year ended December 31,
Industry Averages 2019 2018
Return on equity 5.31% 12.54%
Profit margin 4.00 6.21
Asset turnover 0.83 1.96
Leverage 1.60 1.03

Required:

Perform a Dupont analysis for 2018 and 2019. Round intermediate calculations to four decimals, if required and your final answers to two decimal places.

Return on Equity
2019 %
2018 %
Answer #1

Remington inc.

Return on equity (Dupont analysis)

= Net profit margin ratio × Assets turnover × Financial leverage

= Net income/Sales × Sales/Average total assets × Average total assets/Average shareholders equity

Return on equity %
2019

= (391,179/72,45,088) × {72,45,088/(41,21,005+38,46,450)/2} × {((41,21,005+38,46,450)/2)/(24,27,193+22,67,617)/2}

= 0.0540 × 1.8187 × 1.6971

= 16.67%

2018

= (455,497/69,44,296) × {69,44,296/(38,46,450+34,85,233)/2} × {((38,46,450+34,85,233)/2)/(22,67,617+20,83,122)/2}

= 0.0656 × 1.8943 × 1.6852

= 20.94%